OpenStudy (ac3):

(2pi)Integral(xsqrt(1+(1/(3(x)^(2/3)))^2))

OpenStudy (zale101):

\(\Large 2\pi \int_{}^{} x\sqrt{1+\frac{1}{(3(x)^{2/3})^2}} *dx\) Does the question look like this?

OpenStudy (ac3):